A manufacturer needs to borrow money to purchase a building. The purchase price of the building is $1.5 million, and the company will put $300,000 in cash down at closing. If the company can borrow the difference from its bank at 4.85% for 20 years, what will the monthly principal and interest payment of the loan be? Create an amortization schedule also.
Answer #1

If the down payment = 300000 then the loan amount = 1500000-300000 = 1.2 million

We are given the following information:

r 4.85%
n 20
frequency 12
PV $   12,00,000.00

We need to solve the following equation to arrive at the required monthly payment:

So the total monthly payment is 7820.37 including principal and interest

Amortization schedule is as follows:

Opening balance = previous year's closing balance
Closing balance = Opening balance-Principal repayment
PMT is calculated as per the above formula
Interest = 0.0485 /12 x opening balance
Principal repayment = PMT - Interest